# The Z-Gromov-Wasserstein Distance

### v1

@@ 4.2.3 Path-Connectedness
One of the distinct properties of $\mathcal{M}_{\sim}^{Z,p}$ is that it is always path-connected regardless of $Z$. Intuitively, this is because a $Z$-network is equipped with a measure, which is a continuous object taking values on the real line. This intuition plays a fundamental role in the following proof.
Theorem 40.
$\mathcal{M}_{\sim}^{Z,p}$ is always path-connected.
@@
$X_{t}=(X\amalg Y,\omega_{X}\amalg\omega_{Y},(1-t)\mu_{X}+t\mu_{Y})\quad\mbox{for}\quad t\in[0,1].$
@@
$D^{X}:X\to(X\amalg Y)^{2}\quad\mbox{and}\quad D^{Y}:Y\to(X\amalg Y)^{2}$
@@
$\int_{(X\amalg Y)^{2}}\int_{(X\amalg Y)^{2}}d_{Z}(\omega_{X}\amalg\omega_{Y}(u,u^{\prime}),\omega_{Y}\amalg\omega_{Y}(v,v^{\prime}))^{p}\mu(du\times dv)\nu(du^{\prime}\times dv^{\prime}),$ (65)
where $X$ and $Y$ are abusively being used as shorthand for $D^{X}_{*}\mu_{X}$ and $D^{Y}_{*}\mu_{Y}$, respectively. We note that $I_{XX},I_{YX}$ and $I_{YY}$ vanish since the integrand is identically zero by the definition of $D_{*}^{X}\mu_{X}$ and $D_{*}^{Y}\mu_{Y}$. Therefore,
@@ 4.2.4 Geodesicity
Let $X=X_{0}\times X_{1}$ and construct a coupling of $\pi$ and $\pi$ on $X\times X$ by $\Delta_{*}\pi$ where $\Delta:X\to X\times X$ is the standard diagonal map, $\Delta(x)=(x,x)$. Then, for any $s,t\in[0,1]$,
@@
The argument uses the change of variable formula in (70), that $\omega_{t}$ is a geodesic in (71), the definition of $\omega_{0}$ and $\omega_{1}$ in (72) and that $\pi$ is an optimal coupling in (73). ∎
@@
defines a geodesic between $X_{0}$ and $X_{1}$ because, for $0\leq s\leq t\leq 1$, by Lemma 35, we have $\mathrm{GW}_{1}^{Z}(X_{s},X_{t})\leq\frac{1}{2}D_{1}(\omega_{s},\omega_{t})$, and
@@
Consider $Z$-networks $X=(\{0\},\omega_{0},\delta_{0}),Y=(\{0\},\omega_{1},\delta_{0})$ where $\omega_{0},\omega_{1}$ are the functions with $\omega_{0}(0,0)=0$ and $\omega_{1}(0,0)=1$. Suppose that $\mathcal{M}_{\sim}^{Z,p}$ is geodesic. Then, there is a midpoint between them. [10, Lemma 2.4.8] That is, there exists a $Z$-network $M=(M,\omega_{M},\mu_{M})$ such that
@@
These two results suggest that the geodesic property differs between $p=1$ and $p>1$. In fact, the Wasserstein distance holds such a property: the Wasserstein $p$-space is always geodesic if $p=1$, and it relies on the underlying metric if $p>1$ [54, Theorem 3.16, Remark 3.18]. Considering this result, we conclude this subsection by posing the following:
Questions 44.
  1. 1.
@@ 5.1 Hierarchy of Lower Bounds
Definition 45.
Let $X=(X,\omega_{X},\mu_{X}),(Y,\omega_{Y})$ be a $(Z,p)$-network. We define the following invariants:
@@
    $\textnormal{size}_{p,z_{0}}(X)=\|d_{Z}(\omega_{X}(x,x^{\prime}),z_{0})\|_{L^{p}(\mu_{X}\otimes\mu_{X})}$ (80)
@@
Lemma 48 (Lemma 2.2, [71], Lemma 27, [19]).
@@
Proof of Lemma 47.
The $\subset$ direction is standard [19]. For the opposite direction, we briefly point out that the proof in [19] contains a gap. Let $\pi\in\mathcal{C}(f_{*}\mu_{X},g_{*}\mu_{Y})$, and our goal is to find a measure $\sigma\in\mathcal{C}(\mu_{X},\mu_{Y})$ such that $T_{*}\sigma=\pi$. The strategy in [19] was to apply Lemma 48 directly to the mapping $T$ so that we have a probability measure $\sigma$ such that $\pi=T_{*}\sigma$, but the issue here is that Lemma 48 only ensures that $\sigma$ is a probability measure and not necessarily a coupling. Indeed, if $f,g$ are constant maps, $\mathcal{C}(f_{*}\mu_{X},g_{*}\mu_{Y})$ only consists of a Dirac measure, and importantly, any probability measure $\sigma$ (not necessarily a coupling) on $X\times Y$ satisfies $\pi=T_{*}\sigma$ for the element $\pi\in\mathcal{C}(f_{*}\mu_{X},g_{*}\mu_{Y})$. To impose an additional restriction that $\sigma$ is a coupling, we consider the augmented $T$ mapping $\bar{T}:X\times Y\to X\times Z\times Z\times Y$,
@@
where $\pi_{X}:X\times Y\to X$ is the projection onto the first component and $\pi_{Y}$ similarly. Intuitively, we added the projection functions $\pi_{X},\pi_{Y}$ so that the condition $\mu_{X}\otimes\pi\otimes\mu_{Y}=\bar{T}_{*}\sigma=(\pi_{X},T,\pi_{Y})_{*}\sigma$ we will obtain from Lemma 48 induces $(\pi_{X})_{*}\sigma=\mu_{X},(\pi_{Y})_{*}\sigma=\mu_{Y}$ i.e. the coupling condition.
Since $X,Y,Z$ are Polish, $X\times Y$ and $X\times Z\times Z\times Y$ are again Polish, and $\bar{T}$ can be made surjective by restricting the range to $\bar{T}(X\times Y)=X\times T(X\times Y)\times Y$. The measurability of $f,g$ implies $\bar{T}$ is measurable, so we apply Lemma 48 to the measure $\mu_{X}\otimes\pi|_{T(X\times Y)}\otimes\mu_{Y}$ where $\pi|_{T(X\times Y)}(A)=\pi(A\cap T(X\times Y))$ and the mapping $\bar{T}$ to obtain a probability measure $\sigma$ on $X\times Y$ such that $\mu_{X}\otimes\pi|_{T(X\times Y)}\otimes\mu_{Y}=\bar{T}_{*}\sigma$. $\sigma$ belongs to $\mathcal{C}(\mu_{X},\mu_{Y})$ because, by definition, for any measurable $A\subset X$, we have
@@
i.e., by “hiding” the $\nu$ integral and infimum in the definition of $\textnormal{ecc}_{p,X,Y}^{\textnormal{out}}$. The equality (TLB) $=$ ($Z$-TLB) is obtained by Lemma 47 by setting $f=\omega_{X}(x,\cdot),g=\omega_{Y}(y,\cdot)$. (TLB) $\geq$ (FLB) is obtained by applying the reverse triangle inequality
@@
The equality (FLB) $=$ ($Z$-FLB) is again obtained by Lemma 47 by setting $f=\textnormal{ecc}_{p,z_{0},X}^{\textnormal{out}},g=\textnormal{ecc}_{p,z_{0},Y}^{\textnormal{out}}$. We note that, unlike in ($Z$-TLB), the Wasserstein distance for ($Z$-FLB) is over $\mathbb{R}$. The inequality (FLB) $\geq$ (SzLB) is obtained by another application of the reverse Minkowski inequality,
@@
and the equality (SLB) $=$ ($Z$-SLB) is obtained by Lemma 47 by setting $f=\omega_{X},g=\omega_{Y}$. ∎

### v2

@@ 4.2.3 Path-Connectedness
One of the distinct properties of $\mathcal{M}_{\sim}^{Z,p}$ is that it is always contractible, regardless of the topology of $Z$. We prove this below, but first show that $\mathcal{M}_{\sim}^{Z,p}$ is always path-connected. Intuitively, the result holds because a $Z$-network is equipped with a measure, which is a continuous object taking values on the real line. This intuition plays a fundamental role in the following proof. An important feature of the proof is that it provides an explicit formula for a continuous path between any pair of $Z$-networks. This formula will be used later in the proof of contractibility.
Proposition 4.11.
For any space $Z$, $\mathcal{M}_{\sim}^{Z,p}$ is path-connected for all $p\in[1,\infty)$.
The $p=\infty$ case is more subtle, and is not completely resolved, as we describe below.
@@
$X_{t}=(X\amalg Y,\omega_{X}\amalg\omega_{Y},(1-t)\mu_{X}+t\mu_{Y})\quad\mbox{for}\quad t\in[0,1].$ (70)
@@
$\Delta^{X}:X\to(X\amalg Y)^{2}\quad\mbox{and}\quad\Delta^{Y}:Y\to(X\amalg Y)^{2}$
@@
$\int_{(X\amalg Y)^{2}}\int_{(X\amalg Y)^{2}}d_{Z}(\omega_{X}\amalg\omega_{Y}(u,u^{\prime}),\omega_{X}\amalg\omega_{Y}(v,v^{\prime}))^{p}\mu(du\times dv)\nu(du^{\prime}\times dv^{\prime}),$ (73)
where $X$ and $Y$ are abusively being used as shorthand for $\Delta^{X}_{*}\mu_{X}$ and $\Delta^{Y}_{*}\mu_{Y}$, respectively. We note that $I_{XX},I_{YX}$ and $I_{YY}$ vanish since the integrand is identically zero by the definition of $\Delta_{*}^{X}\mu_{X}$ and $\Delta_{*}^{Y}\mu_{Y}$. Therefore,
@@
Let us now discuss the path-connectivity properties of $\mathcal{M}^{Z,\infty}_{\sim}$. To do so, it will be useful to introduce an invariant of a $Z$-network $X=(X,\omega_{X},\mu_{X})$. For a fixed point $z\in Z$, we define the $p$-size of $X$, relative to $z$ to be
$\mathrm{size}_{p,z}(X)\coloneqq\|d_{Z}(\omega_{X}(\cdot,\cdot),z)\|_{L^{p}(\mu_{X}\otimes\mu_{X})}$ (76)
(see also Definition 5.1 below). It is a fact that, for any $Z$-networks $X$ and $Y$,
$\mathrm{GW}_{p}^{Z}(X,Y)\geq\frac{1}{2}|\mathrm{size}_{p,z}(X)-\mathrm{size}_{p,z}(Y)|.$ (77)
Indeed, we prove this below in Theorem 8, in the context of several lower bounds on the $Z$-GW distance.
We now show that Proposition 4.11 fails in the $p=\infty$ case. Intuitively, the proof strategy does not apply because it is based on manipulating “weights” in the measures, whereas $\mathrm{GW}^{Z}_{\infty}$ is insensitive to these weights and only optimizes a quantity which depends on supports of couplings.
Example 4.12.
Let $Z=\{0,1\}\subset\mathbb{R}$, with the restriction of Euclidean distance. We claim that $\mathcal{M}^{Z,\infty}_{\sim}$ is not path connected. To see this, observe that the size function (76) has the property that
$\mathrm{size}_{\infty,0}(X)\in\{0,1\}$
for all $X\in\mathcal{M}^{Z,\infty}_{\sim}$. By Theorem 8 (or (77)), the function $\mathrm{size}_{\infty,0}:\mathcal{M}^{Z,\infty}_{\sim}\to\mathbb{R}$ is Lipschitz continuous. This means that, if $X_{t}$ is a continuous path in $\mathcal{M}^{Z,\infty}_{\sim}$, then the composition
$t\mapsto\mathrm{size}_{\infty,0}(X_{t}):[0,1]\to\{0,1\}$
must be continuous, and therefore constant. It follows that there is no path joining a $(Z,\infty)$-network of size 00 to one of size $1$. For a specific example, there is no continuous path from
$X=(\{0\},\omega_{X}(0,0)=0,\delta_{0})\quad\mbox{to}\quad Y=(\{1\},\omega_{Y}(1,1)=1,\delta_{1})$
in $\mathcal{M}^{Z,\infty}_{\sim}$.
The idea of the justification in Example 4.12 can be extended to show that path-continuity of $Z$ is a necessary condition for path-continuity of $\mathcal{M}_{\sim}^{Z,\infty}$ (we omit the details of this extension here). However, we were unable to determine whether this condition is sufficient. On the other hand, if $Z$ is geodesic then so is $\mathcal{M}_{\sim}^{Z,\infty}$, as we show below in Theorem 7. This observation illustrates the subtlety of the following question:
Question 1.
Does path-connectivity of $Z$ imply path-connectivity of $\mathcal{M}_{\sim}^{Z,\infty}$?
4.2.4 Contractibility
We now take the path-connectivity result a step further and show that $\mathcal{M}_{\sim}^{Z,p}$ is always contractible.
Theorem 6.
For any space $Z$, $\mathcal{M}_{\sim}^{Z,p}$ is contractible for all $p\in[1,\infty)$.
The proof will use the following lemma, which specializes some features of the proof of Proposition 4.11 to the case where one of the $Z$-networks is a one-point space. The result uses the size functions introduced in (76).
Lemma 4.13.
Let $X=(X,\omega_{X},\mu_{X})$ be an arbitrary $Z$-network, and let $Y=(\{\star\},\omega_{Y},\delta_{\star})$ be a one-point $Z$-network, where $\omega_{Y}(\star,\star)=z$ for some fixed $z\in Z$, and $\delta_{\star}$ denotes the Dirac measure. Let $X_{t}$ denote the path defined in (70) between $X$ and $Y$ (considered up to weak isomorphism), where we specifically define
$\omega_{X}\amalg\omega_{Y}(u,v)=\left\{\begin{array}[]{lr}\omega_{X}(u,v)&\mbox{if }u,v\in X\\ z&\mbox{otherwise.}\end{array}\right.$
Then the Hölder estimate (75) simplifies to
$\mathrm{GW}_{p}^{Z}(X_{s},X_{t})\leq\frac{(3|t-s|)^{1/p}}{2}\cdot\mathrm{size}_{p,z}(X).$ (78)
Proof.
Observe that there is a unique coupling $\pi$ between $\mu_{X}$ and $\delta_{\star}$, characterized by
$\pi(dx\times d\star)=\mu_{X}(dx).$
The integrals $I_{\pi\pi}$, $I_{X\pi}$ and $I_{\pi Y}$ in (75) can then be computed explicitly by applying the definitions:
$I_{\pi\pi}$ $=\iint_{(X\amalg\{\star\})^{4}}d_{Z}(\omega_{X}\amalg\omega_{Y}(u,u^{\prime}),\omega_{X}\amalg\omega_{Y}(v,v^{\prime}))^{p}\pi(du\times dv)\pi(du^{\prime}\times dv^{\prime})$ (79)
$=\iint_{(X\times\{\star\})^{2}}d_{Z}(\omega_{X}\amalg\omega_{Y}(x,x^{\prime}),\omega_{X}\amalg\omega_{Y}(\star,\star))^{p}\pi(dx\times d\star)\pi(dx^{\prime}\times d\star)$ (80)
$=\iint_{X\times X}d_{Z}(\omega_{X}(x,x^{\prime}),z)^{p}\mu_{X}(dx)\mu_{X}(dx^{\prime})=\mathrm{size}_{p,z}(X)^{p},$ (81)
$I_{X\pi}$ $=\iint_{(X\amalg\{\star\})^{4}}d_{Z}(\omega_{X}\amalg\omega_{Y}(u,u^{\prime}),\omega_{X}\amalg\omega_{Y}(v,v^{\prime}))^{p}\Delta^{X}_{\ast}\mu_{X}(du\times dv)\pi(du^{\prime}\times dv^{\prime})$ (82)
$=\int_{X\times\{\star\}}\int_{X}d_{Z}(\omega_{X}\amalg\omega_{Y}(x,x^{\prime}),\omega_{X}\amalg\omega_{Y}(x,\star))^{p}\mu_{X}(dx)\pi(dx^{\prime}\times d\star)$ (83)
$=\iint_{X\times X}d_{Z}(\omega_{X}(x,x^{\prime}),z)^{p}\mu_{X}(dx)\mu_{X}(dx^{\prime})=\mathrm{size}_{p,z}(X)^{p},$ (84)
and, similarly,
$I_{\pi Y}$ $=\iint_{(X\amalg\{\star\})^{4}}d_{Z}(\omega_{X}\amalg\omega_{Y}(u,u^{\prime}),\omega_{X}\amalg\omega_{Y}(v,v^{\prime}))^{p}\pi(du\times dv)\Delta^{Y}_{\ast}\mu_{(}du^{\prime}\times dv^{\prime})$ (85)
$=\int_{X\times\{\star\}}\int_{\{\star\}}d_{Z}(\omega_{X}\amalg\omega_{Y}(x,\star),\omega_{X}\amalg\omega_{Y}(\star,\star))^{p}\pi(dx\times d\star)\delta_{\star}(d\star)$ (86)
$=\int_{X\times\{\star\}}\int_{\{\star\}}d_{Z}(z,z)^{p}\pi(dx\times d\star)=0.$ (87)
The estimate (75) therefore becomes
$2^{p}\mathrm{GW}_{p}^{Z}(X_{s},X_{t})^{p}\leq|t-s|\cdot 3\cdot\mathrm{size}_{p,z}(X)^{p},$
and this can be rewritten as (78). ∎
We now proceed with the proof of Theorem 6. The strategy is to define an explicit contraction to a one-point space which follows paths as defined in the proof of Proposition 4.11. Namely, fix an arbitrary point $z\in Z$ and, for any $Z$-network $X=(X,\omega_{X},\mu_{X})$ and $t\in[0,1]$, let $X_{t}$ denote the $Z$-network
$X_{t}=(X\amalg\{\star\},\hat{\omega}_{X},(1-t)\mu_{X}+t\delta_{\star}),$ (88)
where $\star$ is an abstract point,
$\hat{\omega}_{X}(x,x^{\prime})=\left\{\begin{array}[]{lr}\omega_{X}(x,x^{\prime})&x,x^{\prime}\in X\\ z&\mbox{otherwise,}\end{array}\right.$
and $\delta_{\star}$ is the Dirac measure. This is then the path from $X$ to a one-point space, as in Proposition 4.11 and Lemma 4.13, where we are using the notation $\hat{\omega}_{X}$ rather than $\omega_{X}\amalg\omega_{Y}$ to condense notation. This will be used to construct the contraction.
Proof of Theorem 6.
We define a map $\Phi:\mathcal{M}_{\sim}^{Z,p}\times[0,1]\to\mathcal{M}_{\sim}^{Z,p}$ by
$\Phi\big{(}[X],t)=[X_{t}],$
where we are using $[X]$ to denote the weak isomorphism class of a $Z$-network $X$, and $X_{t}$ is as in (88). Arguments similar to those used in the proof of Proposition 4.11 show that $\Phi$ is well defined, and that, for any $Z$-network $X$, $X_{0}\sim X$, and $X_{1}$ is weakly isomorphic to the $Z$-network
({}$(\{\star\},\omega_{\{\star\}},\delta_{\star}),\quad\mbox{with}\quad\omega_{\{\star\}}(\star,\star)=z,$
so that $\Phi(\cdot,1)$ is a constant map. It remains to show that $\Phi$ is continuous. This will be done in two steps: we show that the component function $\Phi(\cdot,t)$ is Lipschitz continuous for each fixed $t\in[0,1]$, and then combine this with various estimates to derive continuity in general.
Following the plan described above, fix $t\in[0,1]$ and consider the restricted map
$\Phi(\cdot,t):\mathcal{M}_{\sim}^{Z,p}\to\mathcal{M}_{\sim}^{Z,p}.$
Toward establishing its continuity, let $X$ and $Y$ be $Z$-networks and choose an optimal coupling $\pi\in\mathcal{C}(\mu_{X},\mu_{Y})$. Let $X_{t}$ and $Y_{t}$ be as in (88). We extend $\pi$ to a measure on
$W=(X\amalg\{\star\})\times(Y\amalg\{\star\})=(X\times Y)\amalg(X\times\{\star\})\amalg(\{\star\}\times Y)\amalg\{(\star,\star)\}$ (89)
[…]
$2^{p}\mathrm{GW}_{p}^{Z}(X_{t},Y_{t})^{p}$ $\leq\iint_{W\times W}d_{Z}(\hat{\omega}_{X}(x,x^{\prime}),\hat{\omega}_{Y}(y,y^{\prime}))^{p}\overline{\pi}(dx\times dy)\overline{\pi}(dx^{\prime}\times dy^{\prime})$ (90)
$=\left(\iint_{(X\times Y)^{2}}+\iint_{\{(\star,\star)\}^{2}}\right)d_{Z}(\hat{\omega}_{X}(x,x^{\prime}),\hat{\omega}_{Y}(y,y^{\prime}))^{p}\overline{\pi}(dx\times dy)\overline{\pi}(dx^{\prime}\times dy^{\prime})$ (91)
$=(1-t)^{2}\iint_{(X\times Y)^{2}}d_{Z}(\omega_{X}(x,x^{\prime}),\omega_{Y}(y,y^{\prime}))^{p}\pi(dx\times dy)\pi(dx^{\prime}\times dy^{\prime})$ (92)
$=(1-t)^{2}2^{p}\mathrm{GW}_{p}^{Z}(X,Y)^{p},$ (93)
with the various equalities justified as follows: (91) uses the observation above on the support of $\overline{\pi}$, together with the fact that
$d_{Z}(\hat{\omega}_{X}(x,\star),\hat{\omega}_{Y}(y,\star))=d_{Z}(\hat{\omega}_{X}(\star,x),\hat{\omega}_{Y}(\star,y))=0,\qquad\forall\;(x,y)\in X\times Y,$
so that the remaining “cross-term” integrals $\iint_{(X\times Y)\times\{(\star,\star)\}}$ and {(,)}×$\iint_{\{(\star,\star)\}\times(X\times Y)}$ vanish; (92) follows by the definitions of $\hat{\omega}_{X}$, $\hat{\omega}_{Y}$ and $\overline{\pi}$; finally, (93) is given by the optimality of $\pi$. This shows that $\Phi(\cdot,t)$ is Lipschitz continuous.
We now show that the map $\Phi$ is continuous. Fix a $Z$-network $Y$ and $t\in[0,1]$, and let $\epsilon>0$. For the moment, let us assume that $\mathrm{size}_{p,z}(Y)>0$ and $t<1$ (the remaining special case will be addressed later), so that we may assume without loss of generality (for technical reasons) that
$\epsilon<\frac{3}{2}\mathrm{size}_{p,z}(Y)(1-t)^{2/p},\quad\mbox{or}\quad\frac{2\epsilon}{3\mathrm{size}_{p,z}(Y)(1-t)^{2/p}}<1.$ (94)
We will determine $\delta>0$ such that
$\max\{\mathrm{GW}_{p}^{Z}(X,Y),|t-s|\}<\delta\Rightarrow\mathrm{GW}_{p}^{Z}(X_{s},Y_{t})<\epsilon,$
for an arbitrary $Z$-network $X$ and $s\in[0,1]$. We have
$\mathrm{GW}_{p}^{Z}(X_{s},Y_{t})$ $\leq\mathrm{GW}_{p}^{Z}(X_{s},X_{t})+\mathrm{GW}_{p}^{Z}(X_{t},Y_{t})$ (95)
$\leq\frac{(3|t-s|)^{1/p}}{2}\mathrm{size}_{p,z}(X)+(1-t)^{2/p}\mathrm{GW}_{p}^{Z}(X,Y)$ (96)
$\leq\frac{(3|t-s|)^{1/p}}{2}\mathrm{size}_{p,z}(Y)+(3|t-s|)^{1/p}\mathrm{GW}_{p}^{Z}(X,Y)+(1-t)^{2/p}\mathrm{GW}_{p}^{Z}(X,Y)$ (97)
where (96) follows by Lemma 4.13 and the Lipschitz bound (93), and (97) follows by the lower bound (77) on $Z$-GW distance in terms of size functions. Choose $\delta>0$ such that
$\delta<\min\left\{\frac{1}{3}\left(\frac{2\epsilon}{3\mathrm{size}_{p,z}(Y)}\right)^{p},\frac{\epsilon}{3(1-t)^{2/p}}\right\}.$
The assumption $\max\{\mathrm{GW}_{p}^{Z}(X,Y),|t-s|\}<\delta$ implies
$\frac{(3|t-s|)^{1/p}}{2}\mathrm{size}_{p,z}(Y)+(3|t-s|)^{1/p}\mathrm{GW}_{p}^{Z}(X,Y)+(1-t)^{2/p}\mathrm{GW}_{p}^{Z}(X,Y)$ (98)
$<\frac{2\epsilon}{3\mathrm{size}_{p,z}(Y)}\cdot\frac{1}{2}\mathrm{size}_{p,z}(Y)+\frac{2\epsilon}{3\mathrm{size}_{p,z}(Y)}\cdot\frac{\epsilon}{3(1-t)^{2/p}}+(1-t)^{2/p}\cdot\frac{\epsilon}{3(1-t)^{2/p}}$ (99)
$=\frac{\epsilon}{3}+\frac{2\epsilon}{3\mathrm{size}_{p,z}(Y)(1-t)^{2/p}}\cdot\frac{\epsilon}{3}+\frac{\epsilon}{3}<\epsilon,$ (100)
where we have used the assumption (94) to simplify the middle term in (100). This proves continuity of $\Phi$ at any $Z$-network $Y$ with $\mathrm{size}_{p,z}(Y)>0$ and $t<1$.
It remains to consider the cases $\mathrm{size}_{p,z}(Y)=0$ or $t=1$. In either case, $Y_{t}$ is weakly isomorphic to $X_{1}$, and continuity amounts to controlling the size $\mathrm{GW}_{p}^{Z}(X_{s},X_{1})$. Applying Lemma 4.13 and the size bound (77) gives
$\mathrm{GW}_{p}^{Z}(X_{s},Y_{t})=\mathrm{GW}_{p}^{Z}(X_{s},X_{1})$ $\leq\frac{(3(1-s))^{1/p}}{2}\mathrm{size}_{p,z}(X)$ (101)
$\leq\frac{(3(1-s))^{1/p}}{2}(\mathrm{size}_{p,z}(Y)+2\mathrm{GW}_{p}^{Z}(X,Y)).$ (102)
If $t=1$, then (102) can be made arbitrarily small by taking $s$ sufficiently close to $1$ and $X$ sufficiently close to $Y$. If $\mathrm{size}_{p,z}(Y)=0$, then we get
$\frac{(3(1-s))^{1/p}}{2}(\mathrm{size}_{p,z}(Y)+2\mathrm{GW}_{p}^{Z}(X,Y))\leq 3^{1/p}\mathrm{GW}_{p}^{Z}(X,Y),$
which gives control of $\mathrm{GW}_{p}^{Z}(X_{s},Y_{t})$ in terms of $\mathrm{GW}_{p}^{Z}(X,Y)$, and this completes the proof. ∎
In analogy with the discussion in the previous subsection, the proof idea for this result fails in the $p=\infty$ case, which remains open.
Question 2.
How does contractibility of $\mathcal{M}_{\sim}^{Z,\infty}$ depend on that of $Z$?
@@ 4.2.5 Geodesicity
Let $X=X_{0}\times X_{1}$ and construct a coupling of $\pi$ and $\pi$ on $X\times X$ by $\Delta_{*}\pi$ where $\Delta:X\to X\times X$ is the standard diagonal map, $\Delta(x)=(x,x)$. Then, for any $s,t\in[0,1]$, and for $p\in[1,\infty)$,
@@
The argument uses the change of variable formula in (105), that $\omega_{t}$ is a geodesic in (106), the definition of $\omega_{0}$ and $\omega_{1}$ in (107) and that $\pi$ is an optimal coupling in (108). The $p=\infty$ case follows from the observation that the proof can be rewritten in terms of $L^{p}$ norms, in which case it applies directly. ∎
@@
defines a geodesic between $X_{0}$ and $X_{1}$ because, for $0\leq s\leq t\leq 1$, by Lemma 4.7, we have $\mathrm{GW}_{1}^{Z}(X_{s},X_{t})\leq\frac{1}{2}D_{1}(\omega_{s},\omega_{t})$, and
@@
Consider $Z$-networks $X=(\{0\},\omega_{0},\delta_{0}),Y=(\{0\},\omega_{1},\delta_{0})$ where $\omega_{0},\omega_{1}$ are the functions with $\omega_{0}(0,0)=0$ and $\omega_{1}(0,0)=1$. Suppose that $\mathcal{M}_{\sim}^{Z,p}$ is geodesic. Then, there is a midpoint between them. [12, Lemma 2.4.8] That is, there exists a $Z$-network $M=(M,\omega_{M},\mu_{M})$ such that
@@
These two results suggest that the geodesic property differs between $p=1$ and $p>1$. In fact, the Wasserstein distance holds such a property: the Wasserstein $p$-space is always geodesic if $p=1$, and it relies on the underlying metric if $p>1$ [63, Theorem 3.16, Remark 3.18]. Considering this result, we conclude this subsection by posing the following:
Question 3.
@@ 5.1 Hierarchy of Lower Bounds
Definition 5.1.
Let $X=(X,\omega_{X},\mu_{X})$ and $Y=(Y,\omega_{Y},\mu_{Y})$ be $(Z,p)$-networks. We define the following invariants:
@@
    $\textnormal{size}_{p,z_{0}}(X)=\|d_{Z}(\omega_{X}(\cdot,\cdot),z_{0})\|_{L^{p}(\mu_{X}\otimes\mu_{X})}$ (115)
@@
Lemma 5.3 (Lemma 2.2, [83], Lemma 27, [23]).
@@
Proof of Lemma 5.2.
The $\subset$ direction is standard [23]. For the opposite direction, we briefly point out that the proof in [23] contains a gap. Let $\pi\in\mathcal{C}(f_{*}\mu_{X},g_{*}\mu_{Y})$, and our goal is to find a measure $\sigma\in\mathcal{C}(\mu_{X},\mu_{Y})$ such that $T_{*}\sigma=\pi$. The strategy in [23] was to apply Lemma 5.3 directly to the mapping $T$ so that we have a probability measure $\sigma$ such that $\pi=T_{*}\sigma$, but the issue here is that Lemma 5.3 only ensures that $\sigma$ is a probability measure and not necessarily a coupling. Indeed, if $f,g$ are constant maps, $\mathcal{C}(f_{*}\mu_{X},g_{*}\mu_{Y})$ only consists of a Dirac measure, and importantly, any probability measure $\sigma$ (not necessarily a coupling) on $X\times Y$ satisfies $\pi=T_{*}\sigma$ for the element $\pi\in\mathcal{C}(f_{*}\mu_{X},g_{*}\mu_{Y})$. To impose an additional restriction that $\sigma$ is a coupling, we consider the augmented $T$ mapping $\bar{T}:X\times Y\to X\times Z\times Z\times Y$,
@@
where $\pi_{X}:X\times Y\to X$ is the projection onto the first component and $\pi_{Y}$ similarly. Intuitively, we added the projection functions $\pi_{X},\pi_{Y}$ so that the condition $\mu_{X}\otimes\pi\otimes\mu_{Y}=\bar{T}_{*}\sigma=(\pi_{X},T,\pi_{Y})_{*}\sigma$ we will obtain from Lemma 5.3 induces $(\pi_{X})_{*}\sigma=\mu_{X},(\pi_{Y})_{*}\sigma=\mu_{Y}$ i.e. the coupling condition.
Since $X,Y,Z$ are Polish, $X\times Y$ and $X\times Z\times Z\times Y$ are again Polish, and $\bar{T}$ can be made surjective by restricting the range to $\bar{T}(X\times Y)=X\times T(X\times Y)\times Y$. The measurability of $f,g$ implies $\bar{T}$ is measurable, so we apply Lemma 5.3 to the measure $\mu_{X}\otimes\pi|_{T(X\times Y)}\otimes\mu_{Y}$ where $\pi|_{T(X\times Y)}(A)=\pi(A\cap T(X\times Y))$ and the mapping $\bar{T}$ to obtain a probability measure $\sigma$ on $X\times Y$ such that $\mu_{X}\otimes\pi|_{T(X\times Y)}\otimes\mu_{Y}=\bar{T}_{*}\sigma$. $\sigma$ belongs to $\mathcal{C}(\mu_{X},\mu_{Y})$ because, by definition, for any measurable $A\subset X$, we have
@@
i.e., by “hiding” the $\nu$ integral and infimum in the definition of $\textnormal{ecc}_{p,X,Y}^{\textnormal{out}}$. The equality (TLB) $=$ ($Z$-TLB) is obtained by Lemma 5.2 by setting $f=\omega_{X}(x,\cdot),g=\omega_{Y}(y,\cdot)$. (TLB) $\geq$ (FLB) is obtained by applying the reverse triangle inequality
@@
The equality (FLB) $=$ ($Z$-FLB) is again obtained by Lemma 5.2 by setting $f=\textnormal{ecc}_{p,z_{0},X}^{\textnormal{out}},g=\textnormal{ecc}_{p,z_{0},Y}^{\textnormal{out}}$. We note that, unlike in ($Z$-TLB), the Wasserstein distance for ($Z$-FLB) is over $\mathbb{R}$. The inequality (FLB) $\geq$ (SzLB) is obtained by another application of the reverse Minkowski inequality,
@@
and the equality (SLB) $=$ ($Z$-SLB) is obtained by Lemma 5.2 by setting $f=\omega_{X},g=\omega_{Y}$. ∎